Brad is 7 more than 2 times as old as Ben. In 23 years the sum of their ages will be 83. How old is Ben now?
Answer by ewatrrr(24785)


Hi,
Let x and (2x+7)represent Ben's and Brad's age NOW.
In 23 years...***
(x+23) + [(2x+7) + 23] = 83
solving for x
x+ 23 + 2x + 30 = 83
3x + 53 = 83
3x = 30
x = 10, Ben's age NOW. (Brad is now 27)
CHECKING our Answer***
33 + 50 = 83